# Insights Into Neutron Stars From Gravitational Redshifts and Universal Relations

### v1

@@ IV Summary and Conclusion
In this work, we have examined the correlation between various properties of NSs and the gravitational redshift using URs. Our findings indicate that the quantities with the strongest correlation are primarily associated with the $C$ and $\kappa$. The quantities $M\times\bar{f}/\chi$, $\bar{I}$, and $\bar{Q}$ are also found to be universal, with most of the observations lying below the 10% tolerance limit. We further show that $\bar{\lambda}$ and $<c_{s}^{2}>$ are not that universal and also tend to violate the 10% tolerance limit.
@@
Compactness or $C$ is the ratio of the mass and radius of the star. The strong universality between $Z_{g}$, $C$ and $\kappa$ is also a portrayal of the fact that $Z_{g}$ has a direct dependence on the ratio of mass and radius. This correlation also suggests that a higher value of $Z_{g}$ is associated with a higher value in $C$. It shows any future redshift observations with a higher value of $Z_{g}$ will be due to a more compact star. Knowledge about the UR of $M\times\bar{f}/\chi$ can help us estimate the rotational properties if the value of $Z_{g}$ is also known. In the event of an observation of redshift from sources like NS, the ratio $\bar{f}/{\chi}$ can be determined within a certain margin of error, provided the mass of the star is known. For the UR with $\bar{I}$ and $\bar{Q}$ within the observational region of redshifts, the error tolerance lies within the 10% limit. The scenario changes when we analyze the URs with respect to $\bar{\lambda}$ and $<c_{s}^{2}>$. For these two particular parameters, we see a higher deviation in the fitting function, resulting in a violation of the 10% limit. Traditionally, we have seen that the $\bar{I}$, $\bar{\lambda}$, and $\bar{Q}$ show higher universality among themselves (the I-love-Q relations), suggesting that if there exists a universal relation between any two parameters then the third parameter will also be universal. Contrary to previous studies, we see that though $Z_{g}$ is universal with $\bar{I}$ and $\bar{Q}$, it is not universal with $\bar{\lambda}$. From the $<c_{s}^{2}>$-$Z_{g}$ analysis we see that lower values of $<c_{s}^{2}>$ is found for smaller $Z_{g}$ values.
@@
We then utilize the URs and redshift measurements of RBS 1223, RX J0720.4-3125, and RX J1856.5-3754 to offer a theoretical estimate of these parameters using the fitting functions. The theoretical estimates calculated using the fitting functions of URs provide a scope to recover the values of stellar properties that are not directly observable. However, the accuracy of the estimates might vary depending on how universal the relations are. The greater the correlation of a parameter with $Z_{g}$, the higher the accuracy of the theoretical estimate. As the percentage fractional error in UR is less than around $0.1$, predicting the limiting values of $C$ and $\kappa$ from a $Z_{g}$ observation will have minimal deviation in theoretical analysis. Similarly, depending on the universality of a relation, other parameters $\bar{Q}$, $\bar{I}$ and $\bar{\lambda}$ can be depicted with a certain degree of accuracy. For each of the three gravitational redshift measurements, the probable range of each parameter is shown in table 2. Luo et al. [40], using a Bayesian approach, provided mass and radius estimates for three measurements along with 68% confidence intervals for the tidal Love number. While our method differs from theirs, a comparison reveals that our theoretical estimates are in close agreement with their results. As shown in table 3, it is important to note that their results are based on a 68% confidence interval, while our theoretical estimates are derived from a 95% confidence interval based on redshift measurements. We see the $\bar{\lambda}$ estimate of RX J0720.4-3125 closely agrees with the Bayesian estimate. The estimates from RBS 1223 are also well within the estimated standard deviation from their work. The results from RX J1856.5-3754 show large deviations. This is attributed to the fact that this observation is accompanied by significant indeterminacy and lower universality, as seen in the deviation plot of fig. 7. Since the uncertainty in the measured redshift for RX J0720.4-3125 is the least, hence the theoretical estimate closely agrees with the Bayesian estimate. It shows that theoretical estimates are very accurate for observations with less uncertainty and can be used as an alternative for statistical analysis if needed. The entire method outlined in this article may not be the most sophisticated or precise; however, it is certainly one of the feasible ways to extract values of NS properties from redshift observations. In appendix A, we have also shown that the upper limit of $Z_{g}$ for stars obeying the current astrophysical constraints cannot exceed $\leq 0.763$.
@@
Table 3: Comparison of $\bar{\lambda}$ estimates of our work with Luo et al. [40] (for only 4P model with 68% confidence interval).
@@ Acknowledgement
Appendix A Upper limit of $Z_{g}$
Figure 9: Variation of $Z_{g}$ with mass of the stars.
In this section, we provide an upper bound for the gravitational redshift measurement. The mass-radius curves for the EoSs used in this work are shown in fig. 1. As the EoSs follow the astrophysical constraints, they can provide us with an upper limit on the value of $Z_{g}$. In fig. 9, we plot the variation of the mass with the gravitational redshift and find that the maximum value of gravitational redshift attained is: $Z_{g}(max)\leq 0.763$ which further constraints the previous maximum estimates of $\leq 2$ [35, 36, 37]. The range of values of $Z_{g}$ for a $1.4M_{\odot}$ NS can also be seen to agree with the limiting values for the same provided by the Ref [38], i.e. $0.854\geq Z_{g}\geq 0.184$.

### v2

@@ 4 Summary and Conclusion
In this work, we have examined the correlation between various properties of NSs and the gravitational redshift using URs. Our findings indicate that the quantities $M\times\bar{f}/\chi$, $\bar{I}$, and $\bar{Q}$ are quasi-universal, with most of the observations lying below the 10% tolerance limit. We further show that $\bar{\lambda}$ and $c_{s}^{2}>$ tend to violate the 10% tolerance limit in universality.
@@
Knowledge about the UR of $M\times\bar{f}/\chi$ can help us estimate the rotational properties if the value of $Z_{g}$ is also known. In the event of an observation of redshift from sources like NS, the ratio $\bar{f}/{\chi}$ can be determined within a certain margin of error, provided the mass of the star is known. For the UR with $\bar{I}$ and $\bar{Q}$ within the observational region of redshifts, the error tolerance lies within the 10% limit. The scenario changes when we analyze the URs with respect to $\bar{\lambda}$ and $c_{s}^{2}>$. For these two particular parameters, we see a higher deviation in the fitting function, resulting in a violation of the 10% limit. Traditionally, we have seen that the $\bar{I}$, $\bar{\lambda}$, and $\bar{Q}$ show higher universality among themselves (the I-love-Q relations), suggesting that if there exists a universal relation between any two parameters then the third parameter will also be universal. Contrary to previous studies, we see that though $Z_{g}$ is universal with $\bar{I}$ and $\bar{Q}$, it is not universal with $\bar{\lambda}$. From the $c_{s}^{2}>-$$Z_{g}$ analysis we see that lower values of $c_{s}^{2}>$ is found for smaller $Z_{g}$ values.
We then utilize the URs and redshift measurements of RBS 1223, RX J0720.4-3125, and RX J1856.5-3754 to offer a theoretical estimate of these parameters using the fitting functions. The theoretical estimates calculated using the fitting functions of URs provide a scope to recover the values of stellar properties that are not directly observable. However, the accuracy of the estimates might vary depending on how universal the relations are. The greater the correlation of a parameter with $Z_{g}$, the higher the accuracy of the theoretical estimate.
@@
Depending on the universality of a relation, other parameters $\bar{Q}$, $\bar{I}$ and $\bar{\lambda}$ can be depicted with a certain degree of accuracy. For each of the three gravitational redshift measurements, the probable range of each parameter is shown in 2. Luo et al. (Luo_2022, ), using a Bayesian approach, provided mass and radius estimates for three measurements along with 68% confidence intervals for the tidal Love number. While our method differs from theirs, a comparison reveals that our theoretical estimates are in close agreement with their results. As shown in the figure 3, it is important to note that their results are based on a 68% confidence interval, while our theoretical estimates are derived from a 95% confidence interval based on redshift measurements. We see the $\bar{\lambda}$ estimate of RX J0720.4-3125 closely agrees with the Bayesian estimate. The estimates from RBS 1223 are also well within the estimated standard deviation from their work. The results from RX J1856.5-3754 show large deviations. This is attributed to the fact that this observation is accompanied by significant indeterminacy and the relation of $\bar{\lambda}$ with $Z_{g}$ has lower universality, as seen in the deviation plot 5. Since the uncertainty in the measured redshift for RX J0720.4-3125 is the least, the theoretical estimate closely agrees with the Bayesian estimate. Our analysis provides a new window to look into URs with regards to $Z_{g}$ .We have also demonstrated that, for stars consistent with current astrophysical constraints, the upper limit of $Z_{g}$ is $\sim$ $0.763$. It shows that theoretical estimates are accurate for observations with less uncertainty and can be used as an alternative for statistical analysis. The entire method outlined in this article may not be the most sophisticated or precise; however, it is certainly one of the feasible ways to extract values of NS properties from redshift observations. It is to be noted that $Z_{g}$ considered in our work is the surface redshift for a static star. In A we show how our results can change quantitatively upon considering the redshift of the rotating star. We also show that the qualitative analysis will hold true along with the URs.
@@
Table 3: Comparison of $\bar{\lambda}$ estimates of our work with Luo et al. (Luo_2022, ) (for only 4P model with 68% confidence interval).
@@ Acknowledgement
Appendix A Effects of rotation on redshifts
The calculations explored in this study consider the ideal scenario where the surface redshift of static non-rotating stars was used for our calculations. In this section, we calculate the variation of polar redshift (for the rotating stars) with respect to the static redshift for all our EoSs as shown in figure 8. We set up the rotating NS code RNS (stergioulas, ; nozawa, ) (highest grid: DMDIV = 151 and DSDIV = 301 ) with a rotating frequency of $\sim 480$ HzHz. We show the fractional relative difference between the two redshifts ($|\Delta|=|{\frac{\text{Polar}\>Z_{g}-\text{Static}\>Z_{g}}{\text{Static}\>Z_{g}}}|$) in the lower panel of the same figure. As per the figure, the fractional relative difference between them never exceeds $20\%$ and is within the range of our observed redshift measurements. We always find them to be less than $15\%$ (for most EoSs, the difference is less than $10\%$). This shows that if one considers the redshift measurements of the rotating star to compute the universal relations, the qualitative analysis will remain the same, only the coefficients of the fitting parameter may vary.
Figure 8: The upper panel shows the variation of the polar redshift with the static redshift. The lower panel represents the fractional relative difference between them.